OpenStudy (xochital):

x + (x – 8) + 1/2 (x – 8) = 38 what is x? 8 ml 14 ml 18 ml 20 ml

OpenStudy (anonymous):

20 ml

OpenStudy (anonymous):

x + (x – 8) + 1/2 (x – 8) = 38 2x+2(x-8)+(x-8)=2*38 2x+2x-16+x-8=76 5x-24=76 5x=76+24 5x=100 x=100/5 x=20 therefore x =20ml

OpenStudy (xochital):

thank you :)